As an analyst in the valuation team your job is to perform significant financial modeling and analysis. Your company is seeing a new sales strategy that require your input. The strategy will be effective for the upcoming 4 Years. If the company adopts the new strategy, sales will grow at the rate of 15% per year for three years. Other ratios such as: Asset turnover, gross margin, the capital structure and income tax will remain unchanged. However, depreciation would be applicable at 8% of net fixed assets at the starting of the year. Moreover, the target rate of return for the company is 12%. Additional financial information for current year is mentioned below:

 

Income Statement

Sales

50,000

Gross Margin (15%)

7,500

Admin., selling and Distribution expenses (7%)

3,500

Profit before tax

10,000

Tax (35%)

3,500

Profit After Taxes

6,500

Balance Sheet

Fixed Assets

17,000

Current Assets

12,000

Equity

25,000

  


C)    Provide detailed comments should the company proceeds with this new strategy or Not?      (3 Marks)

1
Expert's answer
2021-01-31T19:33:15-0500
"Solution"

This can be solved as shown in the table below.



To get Depreciation, consider the table below.



Value of business year"=\frac{PAT\times100}{Cost\ of\ Equity}"

Value of business year 1"=\frac{6,500\times100}{12}=54,164\\"

Value of business year 2"=\frac{10373\times100}{12}=86,444"

value of business year 3"=\frac{14811\times 100}{12}=123,426"

value of business year 4 "=\frac{19900\times 100}{12}=165,832"


Therefore, value of business is increasing as well as profit, and because of this, the company should proceed with the new strategy.